OpenStudy (anonymous):

What is the 25th term in the following arithmetic sequence? -7, -2, 3, 8,

OpenStudy (welshfella):

what is the common difference of this sequence ? -what is added to one term to get the next?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

right and the first term a1 is -7 the formula for the nth term is a1 + d(n - 1) where a1 = first term (-7) and d = common difference = 5 so for the 25th term plug in these values into the formula ( n = 25)

OpenStudy (anonymous):

127

OpenStudy (anonymous):

?

OpenStudy (welshfella):

-7 + 5(25-1) = -7 + 120

OpenStudy (anonymous):

ohh so 113?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

thank you so much u are a life saver!!!!!!!!!!!!!!

OpenStudy (anonymous):

i have been struggling with this section and i dont know why